### 08 Jan 2024 v1

@@ 5.3 Higgs Vacuum Expectation Value
The Higgs vacuum expectation value $v$ is given by $v_{0}=2\times 10^{-17}\text{m}_{\text{P}}$, where $\text{m}_{\text{P}}$ is the Planck mass. According to [5], its LPI (in units of $m_{\text{P}}$) is given by
@@
If $v$ falls below the lower bound $a$, the proton becomes heavier than the neutron, causing hydrogen to be unstable and enabling electron capture reaction [61]. If $v$ exceeds the upper bound $b$ of $\ell_{v}$, there is no nuclear binding. The Higgs vacuum expectation value is positive as it describes the average value of the Higgs field in the vacuum (i.e., $\mathcal{X}=\mathbb{R}^{+}$). Writing (15) as in (3), we obtain
$\ell_{v}$ $=\frac{a+b}{2}[1-\epsilon,1+\epsilon]$
$=2.04\times 10^{-17}[1-0.6176,1+0.6176].$
Since $\epsilon=0.6176$ is not small, we cannot use (14) in order to calculate the maximum tuning probability as $\text{TP}_{\max}=2e^{-1}\times 0.6176=0.4544$. Instead, we make use of the fact that for an exponential distribution $F(\cdot;\theta)$, with mean $\theta$, the tuning probability of $\ell_{v}$ equals
@@
$\text{TP}_{\text{max}}=\left(\frac{b}{a}\right)^{-a/(b-a)}-\left(\frac{b}{a}\right)^{-b/(b-a)}=0.4886.$ (16)
@@ 5.4 Amplitude of primordial fluctuations
According to Martin [62], the amplitude $Q$ of the primordial fluctuations must be in the interval
@@
in agreement with [21]. Since this number is not small, the FT of $Q$ cannot be known.
@@ 5.5 Baryon-Photon Ratio
According to [4], the baryon-photon ratio $\eta$ has a value $\eta\sim 6\times 10^{-10}$, and it can sustain life in between the interval
@@
Again, since $\epsilon$ is not small, we follow the procedure of the previous two examples in order to calculate the maximal tuning probability
@@
of the ratio $r=b/a=(1+\epsilon)/(1-\epsilon)$ between the right and left end points of $\ell_{X}$. It can be seen that $h:(1,\infty)\to(0,1)$ is a strictly increasing function of $r$, with $\lim_{r\to 1}h(r)=:h(1)=0$, $h^{\prime}(1)=e^{-1}$ and $\lim_{r\to\infty}h(r)=1$. From this, it follows that (14) is a special case of (19) which holds in the limit of small half relative sizes $\epsilon>0$ of $\ell_{X}$. Moreover, it follows from Theorem 3 that a sufficient condition to know about FT of $X$ on $\mathcal{X}=\mathbb{R}^{+}$ is
@@ 6 Conclusion
Using the recent mathematization of knowledge acquisition in [42], we formalize Principle 1, which establishes that the FT of a given physical parameter $X$ can be known at least if the following two conditions are satisfied
@@
    The constant $X$ can only take non-negative values.
@@
    The size of the life-permitting interval (LPI) is small relative to the observed value $x_{0}$ of the constant in our universe.
This latter proviso is extremely important, since as we saw in the examples, it is even possible to have very small intervals whose size relative to $x_{0}$ is still large. As a result, (B) is not satisfied and the level of tuning cannot be known. In other words, it is the relative half size of the interval that must be small, not the interval itself, to know that a given constant is fine-tuned. That is, $|\mbox{LPI}|\ll x_{0}$ is required in order to satisfy (B). Our conclusion is sobering in the sense that many things that, in spite of the ado, are touted as fine-tuned cannot be known to be so.

### 01 Mar 2024 v2

@@ 5.3 Higgs Vacuum Expectation Value
The Higgs vacuum expectation value $v$ is given by $v_{0}=2\times 10^{-17}\text{m}_{\text{P}}$, where $\text{m}_{\text{P}}$ is the Planck mass. According to Barnes (2012), its LPI (in units of $m_{\text{P}}$) is given by
@@
If $v$ falls below the lower bound $a$, the proton becomes heavier than the neutron, causing hydrogen to be unstable and enabling electron capture reaction (Damour & Donoghue, 2008). If $v$ exceeds the upper bound $b$ of $\ell_{v}$, there is no nuclear binding. The Higgs vacuum expectation value is positive as it describes the average value of the Higgs field in the vacuum (i.e., $\mathcal{X}=\mathbb{R}^{+}$). Writing (15) as in (3), we obtain
$\ell_{v}=\frac{a+b}{2}[1-\epsilon,1+\epsilon]=2.04\times 10^{-17}[1-0.617,1+0.617].$
Since $\epsilon=0.617$ is not small, we cannot use (14) to calculate the maximum tuning probability as $\text{TP}_{\max}=2e^{-1}\times 0.617=0.454$. Instead, we make use of the fact that for an exponential distribution $F(\cdot;\theta)$, with mean $\theta$, the tuning probability of $\ell_{v}$ equals
@@
$\text{TP}_{\text{max}}=\left(\frac{b}{a}\right)^{-a/(b-a)}-\left(\frac{b}{a}\right)^{-b/(b-a)}=0.488.$ (16)
@@ 5.4 Amplitude of primordial fluctuations
According to Rees (2000), the LPI of the amplitude $Q$ of the primordial fluctuations is
@@
in agreement with Díaz-Pachón et al. (2021). Since this number is not small, the FT of $Q$ cannot be known.
@@ 5.5 Baryon-Photon Ratio
According to Adams (2019), the baryon-photon ratio $\eta$ has a value $\eta\sim 6\times 10^{-10}$, and it can sustain life in between the interval
@@
Again, since $\epsilon$ is not small, we follow the procedure of the previous two examples to calculate the maximal tuning probability
@@
of the ratio $r=b/a=(1+\epsilon)/(1-\epsilon)$ between the right and left end points of $\ell_{X}$. It can be seen that $h:(1,\infty)\to(0,1)$ is a strictly increasing function of $r$, with $\lim_{r\to 1}h(r)=:h(1)=0$, $h^{\prime}(1)=e^{-1}$ and $\lim_{r\to\infty}h(r)=1$. From this, it follows that (14) is a special case of (19), which holds in the limit of small half relative sizes $\epsilon>0$ of $\ell_{X}$. Moreover, from Theorem 3, a sufficient condition to know about FT of $X$ on $\mathcal{X}=\mathbb{R}^{+}$ is
@@ 6 Conclusion
Using the recent mathematization of knowledge acquisition in (Hössjer et al., 2022), we formalize Principle 1, which establishes that the FT of a given physical parameter $X$ can be known at least if the following two conditions are satisfied:
@@
    The constant $X$ can only take nonnegative values.
@@
    The size of its life-permitting interval is small relative to the observed value $x_{0}$ of the constant in our universe.
This latter proviso is extremely important since, as we saw in the examples, it is possible to have very small intervals whose size relative to $x_{0}$ is large. As a result, (B) is not satisfied and the level of tuning cannot be known. In other words, to know that a given constant is fine-tuned, it is the relative half size of the interval that must be small, not the interval itself. That is, $|\ell_{X}|\ll x_{0}$ is required in order to satisfy (B). Our conclusion is sobering in the sense that many things that, in spite of the ado, are touted as fine-tuned cannot be known to be so.